If the centripetal acceleration for an object is 10m/s2 , Calculate the centripetal acceleration for the same object if its velocity is doubled and radius of rotation decreases to its half value
1
Expert's answer
2019-02-20T13:56:08-0500

The centripetal acceleration

a=v^2/R

Therefore

a^'=v^'2/R'=(2v)^2/(R/2)=8 v^2/R=8a=8×10 m/s^2 =80 m/s^2